This problem seems very simple but I have been stuck on it for 20 minutes and I don't know why!
x^2-x=0
The solution given is x=(0,1) but I do not understand how it was derived. Thankyou!
First you need to look how can you break this down.
so x^2-x = 0, well if you factor out x you have x(x-1)= 0. That means either that x = 0 or x-1=0 solve for x and x = 0 or 1

Found 2 solutions by checkley71, jecegc:
Answer by checkley71(8403)

X^2-X=0
X(X-1)=0
X=0 ANSWER.
X-1=0
X=1 ANSWER.

Answer by jecegc(3)

First you need to look how can you break this down.
so x^2-x = 0, well if you factor out x you have x(x-1)= 0. That means either that x = 0 or x-1=0 solve for x and x = 0 or 1